Solve each of the following systems by substitution.
5x - 2y = -5
y - 5x = 3

Answer by praseenakos@yahoo.com(507)

QUESTION:

Solve each of the following systems by substitution.
5x - 2y = -5
y - 5x = 3


ANSWER:


5x - 2y = -5 ------------(1)

y - 5x = 3 ---------------(2)

Add 5x to both sides of the equation (2)


==> y - 5x + 5x = 3 + 5x


==> y = 3 + 5x --------------(3)


Substitute this value of y, that is y = 3 + 5x in equation(1)


==> 5x - 2(3 + 5x) = -5


==> 5x - 2*3 - 2*5x = -5


==> 5x - 6 - 10x = -5


==> 5x - 10x - 6 = -5

==> -5x - 6 = -5

Add 6 to both sides of the equation....

==> -5x - 6 + 6 = -5 + 6


==> -5x = 1


Divide both sides by -5


==> -5x/-5 = 1/-5



==> x = -1/5


Now put the value of x in (3)


==> y = 3 + 5(-1/5)


==> y = 3 + (-1)


==> y = 2



So we have ,

x = -1/5 and
y = 2



To check your answer, substitute these values in the first equation then you will find it satisfies the equation.


Hope you found the explanation useful...

Regards.
Praseena.